OpenStudy (anonymous):

If \(\int_{2}^{7} (6 f(x) + 7) \, dx = 11\), then \(\int_{2}^{7} f(x) dx =\)

OpenStudy (anonymous):

nobody knows?

OpenStudy (anonymous):

i am trying @Jgeurts

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

my first answer came out to 11/6 but that was wrong

OpenStudy (anonymous):

studying for a midterm so an answer i can reverse engineer into understanding it would be fantastic :p

ganeshie8 (ganeshie8):

\(\large \int \limits_{2}^{7} (6 f(x) + 7) \, dx = 11\) \(\large \int \limits_{2}^{7} 6 f(x) dx + \int \limits_{2}^{7} 7 \, dx = 11\) \(\large 6\int \limits_{2}^{7} f(x) dx + 7(7-2) = 11\) \(\large 6\int \limits_{2}^{7} f(x) dx = -24 \) \(\large \int \limits_{2}^{7} f(x) dx = ?\)

OpenStudy (anonymous):

suddenly, its so easy, thank you so much, -4

ganeshie8 (ganeshie8):

np :)